# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v2

@@ Abstract
Post-hoc, unsupervised concept-based explanation methods (U-CBEMs) translate a vision model’s internal reasoning into human-understandable concepts, leading to interpretable explanations. However, we find that many state-of-the-art (SOTA) U-CBEMs are not faithful: their concepts seem interpretable but fail to reproduce the model’s predictions. We argue that this deficiency has gone unnoticed due to fragmented evaluation—each paper proposes its own faithfulness measure, with no measure-over-measure comparison or broad benchmarking. We close this gap by (i) organizing prior metrics in a unified framework, discussing their limitations, and identifying desiderata for a faithfulness measure; (ii) introducing the Surrogate Faithfulness (SURF) measure, which quantifies faithfulness via the predictive loss of a surrogate that maps explanations to the model’s outputs; and (iii) delivering the first comprehensive U-CBEM faithfulness benchmark across diverse tasks and architectures. In a controlled setting, SURF outperforms prior faithfulness measures in measure-over-measure comparisons, and applying SURF to SOTA U-CBEMs reveals that many visually appealing U-CBEMs are surprisingly unfaithful. We demonstrate SURF applicability in two downstream settings—(i) faithfulness versus the number of concepts used in the explanation and (ii) U-CBEM robustness to adversarial attacks—underscoring SURF’s value as a reliable faithfulness measure. Code to be released.
@@ 1 Introduction
Deep learning models have delivered state-of-the-art results across diverse tasks, yet their internal reasoning remain difficult to interpret [12], which is especially problematic in high-stakes domains such as healthcare and finance. This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior. Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that accurately reflect the model’s internal reasoning and allow humans to reliably predict the model’s output. XAI methods must balance the natural tension between interpretability and faithfulness. While interpretability is typically assessed with human studies, faithfulness can be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model. Because explanations are inevitably lossy, some discrepancy between the surrogate and the model is expected; a faithfulness measure must therefore specify both (i) an appropriate surrogate and (ii) a metric for the induced loss.
@@
One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model. Since the explanation and the model are merged, these approaches eliminate the need for faithfulness checks; however, they often underperform black-box models by reducing model complexity (capacity) for explainability. In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution—commonly found through gradients or perturbations [38, 30, 23, 28, 7]. For images, these attributions are visualized as pixel-level heatmaps. Although such explanations are intuitive, attributions are usually unfaithful, since they are estimated from noisy gradients in deep neural networks or from unrealistic perturbations on the input space [1, 11, 32, 15]. The heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [26, 4].
@@
Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [10, 3, 17, 46]. To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss aspects of the model’s reasoning, raising faithfulness concerns [29]. To avoid annotation, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output.
We argue that, in practice, U-CBEMs have prioritized interpretability at the expense of faithfulness, yielding explanations that look compelling but are surprisingly unfaithful. Despite the importance of faithful explanations, we find that all prior works introduce new faithfulness measures, with no measure-over-measure comparison and sparse cross-U-CBEM evaluations. Thus, there is a lack of consensus on faithfulness measures for U-CBEMs, limiting our understanding of how faithful current approaches are.
This work serves to address this gap. Our contributions are to:
@@
    Organize prior U-CBEM faithfulness measures under a common framework, allowing us to discuss limitations in current measures and identify appropriate desidarata for a faithfulness measure.
@@
    Propose Surrogate Faithfulness (SURF), a faithfulness measure satisfying the desidarata. We conduct a measure-over-measure comparison to empirically validate our faithfulness measure against prior measures.
@@
    Conduct the first, comprehensive faithfulness benchmark of current U-CBEMs across a variety of tasks and model architectures. Most SOTA U-CBEMs that were thought to be faithful are found not to be.
@@
    Extend SURF to two downstream applications. First, studying faithfulness as a function of the number of concepts shows that some U-CBEMs do not get more faithful as they discover more concepts. Second, analyzing the faithfulness of U-CBEMs when the model is under adversarial attack reveals robustness issues.
@@ 2 Preliminaries
We first introduce notation that will be used in the remainder of the paper, as well as preliminaries to familiarize the reader with the field.
Background: Model $\phi:\mathcal{X}\to\mathcal{Y}$ maps from an input space $\mathcal{X}\subseteq\mathbb{R}^{P}$ to an output space $\mathcal{Y}\subseteq\mathbb{R}^{C}$. Assume $\phi$ admits an intermediate space $\mathcal{H}\subseteq\mathbb{R}^{D}$. Let $g:\mathcal{X}\to\mathcal{H}$ and $f:\mathcal{H}\to\mathcal{Y}$. Thus, $\phi(\mathbf{X})=f(g(\mathbf{X}))$. Let $g(\mathbf{x})=\mathbf{h}\in\mathcal{H}$ represent the embedding of $\mathbf{x}$ (e.g., image patch), and $g(\mathbf{X})=\mathbf{H}\in\mathbb{R}^{(HW)\times D}$ denotes the vectorized application of $g$ on $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{HW}\end{bmatrix}^{T}$, where $H,W$ are the spatial dimensions.
Let $E$ be an explanation function, and let $\mathbf{e}=E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image). Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate. Motivated by Ribeiro et al. [30], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$. Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
@@
U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and their concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$. Note that while $K$ can vary per output, all prior works find the same number of CAVs for each output. U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$. The explanation for output $i$ is:
@@
Following prior works, we study faithfulness in the final layer, where the spatial dimensions $H=W=1$, so $\mathbf{H}=\mathbf{h}$. Given that U-CBEMs operate on $g(X)$, the faithfulness measure can be simplified to:
@@ 3.1 Unsupervised Concept-Based Explanations
Wang et al. [37] initiated this line of work, finding CAVs by K-Means clustering intermediate embeddings computed on a dataset. Each cluster was associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches. Ghorbani et al. [13] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate. To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al. [39] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [41] to find concept importance. ICE [44] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation. The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA). Building on this, Fel et al. [9] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance. CDISCO [14] uses the singular value decomposition (SVD) to find CAVs and a gradient-based method for finding concept importance. Recently, MCD [35] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
Table 1: Deletion-Based Measures: We organize deletion-based measures introduced by prior works according to their differing factors. The factors we use are Perturbation Space, Deletion Strategy, and how the results are presented (Y-axis and X-axis). Notice how all prior works introduce a different faithfulness measure, with no measure-over-measure comparison.
Method Perturbation Space Deletion Strategy Y-axis X-axis
@@
ACE Pixel Constant (grey) Accuracy # Deleted Concepts
@@
CDISCO (M1) Pixel Constant (grey) # Classes with $80\%$ accuracy loss # Deleted Concepts
MCD Pixel Inpainting Accuracy Fraction of Inpainted Pixels
CRAFT Concept Constant (zero) Change in Class Logit Score # Deleted Concepts
@@
CDISCO (M2) Weight Constant (zero) Accuracy # Deleted Concepts
3.2 Concept Bottleneck Models
@@
U-CBEMs are not to be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model. These methods [20, 18, 42, 27] train a model to first predict a set of interpretable concepts (e.g., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification). By doing so, CBMs fully decomposes the model’s prediction into interpretable concepts, and exposing the concept representation allows users to intervene (e.g., increase the strength of the “red” concept to help the model recognize the “stop sign”). While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts. As CBMs are inherently explainable, the notion of faithfulness does not extend to them. This work will only discuss faithfulness as it pertains to U-CBEMs.
@@ 4 Unifying U-CBEM Faithfulness Measures
We provide a comprehensive review of all faithfulness measures that were introduced by the works mentioned in Sec. 3.1. We find that each work uses a different measure for faithfulness, indicating a lack of unification in the field and preventing fair comparisons of faithfulness across different U-CBEMs. As part of our unified framework, we place prior faithfulness measures into two categories: 1) deletion-based and 2) surrogate-based. We detail the dimensions across which prior deletion-based measures differ; this is followed by a discussion on their limitations, which raise questions on the validity of past assessments. Then, we similarly characterize prior surrogate-based measures and demonstrate their drawbacks.
4.1 Deletion-Based Measures
Deletion-based measures evaluate faithfulness by observing the effect of deleting a concept on the model’s output. Intuitively, these measures assert that removing the most important concepts will result in the greatest model degradation. After obtaining CAVs and their importances (for a class), concepts are removed (Deletion Strategy) from the input’s Perturbation Space in decreasing order of importance, and the effect on the model is measured after each deletion. Results are visualized by plotting the effect on the model (Y-axis) with respect to some measure of how many concepts were deleted (X-axis). Prior works make different choices along each italicized dimension mentioned.
Perturbation Space: defines the space in which concepts are removed. This is either the image space, the model’s weight space, or the concept space. If making removals in the pixel or the weight space, the U-CBEM must have some mechanism to associate the concept to this space (e.g., which pixels in the image represent the concept).

### v3

@@ Abstract
Deep vision models perform input-output computations that are hard to interpret. Concept-based explanation methods (CBEMs) increase interpretability by re-expressing parts of the model with human-understandable semantic units, or concepts. Checking if the derived explanations are faithful—that is, they represent the model’s internal computation—requires a surrogate that combines concepts to compute the output. Simplifications made for interpretability inevitably reduce faithfulness, resulting in a tradeoff between the two. State-of-the-art unsupervised CBEMs (U-CBEMs) have reported increasingly interpretable concepts, while also being more faithful to the model. However, we observe that the reported improvement in faithfulness artificially results from either (1) using overly complex surrogates, which introduces an unmeasured cost to the explanation’s interpretability, or (2) relying on deletion-based approaches that, as we demonstrate, do not properly measure faithfulness. We propose Surrogate Faithfulness (SURF), which (1) replaces prior complex surrogates with a simple, linear surrogate that measures faithfulness without changing the explanation’s interpretability and (2) introduces well-motivated metrics that assess loss across all output classes, not just the predicted class. We validate SURF with a measure-over-measure study by proposing a simple sanity check—explanations with random concepts should be less faithful—which prior surrogates fail. SURF enables the first reliable faithfulness benchmark of U-CBEMs, revealing that many visually compelling U-CBEMs are not faithful. Code to be released.
@@ 1 Introduction
Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [Gerlings2020ReviewingTN], which is especially problematic in high-stakes domains such as healthcare and finance. This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior. Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that reflect the model’s internal computation. XAI methods must balance the natural tension between interpretability and faithfulness. While interpretability is assessed with human studies, faithfulness must be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model. Since explanations are inevitably lossy, some discrepancy is expected. For certain XAI methods, defining a surrogate is challenging, so alternative measures, or proxies, attempt to evaluate faithfulness using different criteria.
@@
One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [protopnet, senn]. Merging the explanation and the model eliminates the need for faithfulness checks; however, these methods often underperform black-box models by reducing model complexity (capacity) for explainability. In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [wang2024gradient, lime, shapley, Petsiuk2018rise, sobol]. For images, attributions are visualized as pixel-level heatmaps. Although such explanations are intuitive, developing faithfulness measures has been challenging, since many attribution methods do not specify a surrogate to reconstruct the model’s output from the explanation. Thus, proxies—such as checking if important pixels come from foreground regions or looking for significant model degradation after deleting important pixels—are used to study faithfulness. Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [Adebayo2018SanityCF, Frye2020ShapleyManifold, Sixt2019WhenEL, Hase2021TheOP]. The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [Nguyen2021TheEO, colin2022what].
@@
Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [net2vec, netdiss, kimTCAV, zhou2014object]. To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [Ramaswamy2022OverlookedFI]. To avoid these issues, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output. Since CBEMs operate on the model’s intermediate representations, defining a surrogate is more natural, paving the way for proper faithfulness evaluations.
However, a closer look shows that U-CBEMs have adopted inadequate faithfulness measures. We discover that existing evaluations rely on complex surrogates, which allow U-CBEMs to simultaneously show users interpretable explanations and report high faithfulness, yet the explanations do not clearly lead to the model’s output. Furthermore, U-CBEMs adapt popular deletion-based proxies from the attribution literature, yet Sec. 4.1 outlines serious unresolved limitations that prevent them from properly measuring faithfulness. Finally, each work that proposes a new U-CBEM also evaluates it with a new faithfulness measure, with no measure-over-measure comparison, indicating a lack of consensus on faithfulness measures in the field.
We argue that these issues have misled us into believing that current U-CBEMs are faithful. To shed light on this, we make the following contributions:
@@
    Organize prior U-CBEM faithfulness measures under a common framework, allowing us to discuss their limitations and identify appropriate desiderata.
@@
    Propose Surrogate Faithfulness (SURF), a faithfulness measure satisfying the desiderata. Our measure-over-measure comparison checks if faithfulness decreases as the explanation is increasingly randomized; only SURF passes this check.
@@
    Conduct the first, comprehensive faithfulness benchmark of current U-CBEMs across a variety of tasks and model architectures. SURF shows that SOTA U-CBEMs previously evaluated to be faithful are not.
@@
    Leverage SURF to provide a selection criterion for the number of concepts U-CBEMs discover, improving on prior work that sets this hyperparameter arbitrarily.
@@ 2 Preliminaries
We first introduce notation used throughout the paper, as well as preliminaries to familiarize the reader with the field.
Background: Model $\phi:\mathcal{X}\to\mathcal{Y}$ maps from an input space $\mathcal{X}\subseteq\mathbb{R}^{P}$ to an output space $\mathcal{Y}\subseteq\mathbb{R}^{C}$. Assume $\phi$ admits an intermediate space $\mathcal{H}\subseteq\mathbb{R}^{D}$. Let $g:\mathcal{X}\to\mathcal{H}$ and $f:\mathcal{H}\to\mathcal{Y}$. Thus, $\mathbf{y}=\phi(\mathbf{X})=f(g(\mathbf{X}))$. Let $g(\mathbf{x})=\mathbf{h}\in\mathcal{H}$ represent the embedding of $\mathbf{x}$ (e.g., image patch), and $g(\mathbf{X})=\mathbf{H}\in\mathbb{R}^{(HW)\times D}$ denote the vectorized application of $g$ on $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{HW}\end{bmatrix}^{T}$, where $H,W$ are the spatial dimensions.
Let $E$ be an explanation function, and let $E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image). Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate. Motivated by Ribeiro et al. [lime], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$. Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
@@
U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$. Note that while $K$ can vary per output, all prior works find the same $K$ per output. U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$. The explanation for output $i$ is:
@@
Following prior works, we study faithfulness in the final layer, where the spatial dimensions $H=W=1$, so $\mathbf{H}=\mathbf{h}$. Given that U-CBEMs operate on $g(\mathbf{X})$, the faithfulness measure can be simplified to:
@@
Figure 1: Framework. U-CBEM faithfulness measures compare (using metric $d$) the output $\mathbf{y}$ of the original model $\phi(.)=f(g(.))$ with the output $\mathbf{\hat{y}}$ from the explanation. To obtain the explanation, U-CBEMs transform intermediate representation $\mathbf{h}\in\mathcal{R}^{D}$ to the concept representation through concept projection $\mathcal{P}(\mathbf{h})$ and provide an accompanying concept importance AA. The explanation is then passed through surrogate $s$ to obtain $\mathbf{\hat{y}}$. Deletion-based proxies (right) observe model degradation after performing deletion in a deletion space. Surrogate-based measures (left) do not manipulate the computation; instead, they introduce a surrogate to directly approximate Eq. 3.
@@ 3 Related Works
Unsupervised Concept-Based Explanation Methods. Wang et al. [Wang2015UnsupervisedLO] find CAVs by K-Means clustering intermediate embeddings computed on a dataset. Each cluster is associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches. Ghorbani et al. [Ghorbani2019ACE] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate. To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al. [ConceptSHAP] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [Young1985MonotonicSO] to find concept importance. ICE [Zhang2020InvertibleCE] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation. The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA). Building on this, Fel et al. [Fel2022CRAFTCR] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance. CDISCO [graziani2023concept] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance. MCD [vielhaben2023multidimensional] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts. Finding issues with MCD’s interpretability, HU-MCD [hu-mcd] incorporates Segment Anything Model [SAM] to discover more interpretable concepts.
Table 1: Prior deletion-based proxies (top) & surrogate-based measures (bottom) are organized according to their differing factors: Deletion Space (DS), Deletion Method (DM), Metric, Surrogate, and Concept Projection (CP). Each prior work uses a different proxy or measure, with no measure-over-measure comparison.
Method DS DM Metric Surrogate CP
@@
ACE Pixel Constant (grey) Class Accuracy Original Model Identity
MCD Pixel Inpainting Class Accuracy Original Model Identity
HU-MCD Pixel Masking Class Accuracy Original Model Identity
@@
CDISCO (M1) Pixel Constant (grey) # Classes with $80\%$ Accuracy Loss Original Model Identity
@@
CDISCO (M2) Weight Constant (zero) Class Accuracy Original Model Identity
CRAFT Concept Constant (zero) Class Logit Reconstruct $\to$ Original Model U-CBEM
ICE Normalized Target Class L1 Logit Loss U-CBEM Reconstruct $\to$ Original Model U-CBEM
Top-1 Accuracy
C-SHAP Normalized Top-1 Accuracy MLP Reconstruct $\to$ Original Model U-CBEM
@@
Concept Bottleneck Models. CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model. These methods [pmlr-cbm-koh20a, prob-cbm, posthoc-cbm, labelfree-cbm] train a model to first predict a set of interpretable concepts (i.e., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification). By doing so, CBMs fully decompose the prediction into interpretable concepts. While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts. As CBMs are inherently explainable, the notion of faithfulness does not extend to them. This work discusses faithfulness only for U-CBEMs.
@@ 4 Unifying U-CBEM Faithfulness Measures
Each U-CBEM mentioned in Sec. 3 introduces a different faithfulness measure, indicating a lack of consensus and preventing fair comparisons of faithfulness across U-CBEMs. We unify faithfulness measures under a common framework, shown in Fig. 1. Any evaluation measure compares the output of the original model with the output from the explanation (Metric). To obtain the output from the explanation, a Concept Projection transforms the intermediate representation to a concept representation, which is then passed through a Surrogate. Deletion-based proxies assert that removing the most important concepts will result in the greatest model degradation; more degradation is interpreted as a signal for higher faithfulness. Alternatively, surrogate-based measures avoid changing the input and introduce a surrogate to directly approximate faithfulness as in Eq. 3.
4.1 Deletion-Based Proxies
To perform deletion, deletion-based proxies must remove concepts (Deletion Method) from the input’s Deletion Space in decreasing order of importance.
Deletion Space: Where concepts are removed. This is either the image space, the model’s weight space, or the concept space (shown in Fig. 1 (right)).